OpenStudy (anonymous):

Determine which line is parallel to 6x + 6y = 6 1. 2y − 3 = 6x 2. 6y − 3 = x 3. 6x = y + 3 4. None of them. 5. 3x − 6 = −3

OpenStudy (bahrom7893):

parallel = slopes are the same

OpenStudy (bahrom7893):

6x+6y=6 x+y=1 y=-x+1 So slope of ur new line has to be -1

OpenStudy (bahrom7893):

Check the first one: 2y-3=6x; y=3x-3/2.. nope slope's different

OpenStudy (bahrom7893):

repeat the process for each option